Solve
sqrt(x) +5=sqrt(x+55)

Answer by stanbon(75887)

Solve
sqrt(x) +5=sqrt(x+55)
---
Sqruare both sides to get:
x+ 10sqrt(x) + 25 = x+55
---
10sqrt(x) = 30
---
sqrt(x) = 3
----
x = 9
--------------
Checking:
sqrt(9) + 5 = sqrt(9+55)
3 + 5 = 8
==================
Cheers,
Stan H.
==================